OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. -12x + 27y = 60 -6x – 6y = 30

OpenStudy (anonymous):

a. -12x + 27y = 60 b. -6x – 6y = 30 a. (1) -12x +27y=60 b. (-2) 12x +6y=-60 33y=0 y=-33 -6x-6(-33)=60 -6x -198=60 6x=258 x=43 (43,-33)

OpenStudy (anonymous):

You have the right idea when you multiply the second equation by -2 but you need to multiply each term so it should be 12x+12y=-60 then when you add equation 1 and modified equaiton 2 you get 39y=0 and the answer to this is not y =39 but is y =0 (divide both sides by 39). Then substitutie y =0 into either eqaution 1 or two and solve for x (I think youget x=-5